OpenStudy (anonymous):

cos(Sin^(-1)1/2) Without calculator..

OpenStudy (anonymous):

\[\cos (\sin^{-1} 1/2)\]

OpenStudy (lgbasallote):

hit: \[\sin 30 = \frac 12\] \[\sin 60 = \frac {\sqrt 3}{2}\] \[\sin 45 = \frac{\sqrt 2}{2}\] \[\sin 0 = 0\] \[\sin 90 = 1\] so what do you think is \[\sin^{-1} (1/2)\]

OpenStudy (anonymous):